OpenStudy (anonymous):

Use the limit definition to find the slope of the tangent line to the graph of f at the given point. f(x) =Squareroot of x + 1; (3, 2). I got 1/6 but it is incorrect for some reason. Can someone please help?

OpenStudy (precal):

what are you using for the limit definition?

OpenStudy (larseighner):

roots are generally easier to work with when expressed as exponents, so I assume the function is f(x) = (x+1)^(1/2) lim h->0 [f(x+h) - f(x)]/h = lim h=>0 {[(3 + h) + 1]^(1/2) - 2}/h = lim h=>0 {[(3 + h) + 1]^(1/2) - 2}{[(3 + h) + 1]^(1/2) + 2}/(h{[(3 + h) + 1]^(1/2) + 2}) in other words we multiplied both the numerator and denominator by [(3 + h) + 1]^(1/2) + 2, which is equivalent to multiply by 1 and changes nothing. = lim h->0 {[(3 + h) + 1]^(1/2) - 2}^2/(h{[(3 + h) + 1]^(1/2) + 2}) = lim h->0 [(3 + h + 1) - 4]/(h{[(3 + h) + 1]^(1/2) + 2}) = lim h->0 h/(h{[(3 + h) + 1]^(1/2) + 2}) Lo and behold, the h's cancel then the limit can be solved by substitution = lim h->0 1/{[(3 + h) + 1]^(1/2) + 2} = 1/[(4)^(1/2) +2] = 1/4

OpenStudy (anonymous):

From the video I watched to follow the steps they kept it as square root until the end and they squared the number.. But I think the limit definition he used was the limit delta x as it approaches 0

ganeshie8 (ganeshie8):

still looking for help ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

okay :) may i know where exactly are u stuck

OpenStudy (anonymous):

I get 1/6

OpenStudy (anonymous):

I'm not sure where I go worng. I followed all the same steps the video did

ganeshie8 (ganeshie8):

\[\large f'(x) = \lim \limits_{\Delta x \to 0}~ \dfrac{f(x+\Delta x) - f(x)}{\Delta x}\]

ganeshie8 (ganeshie8):

using that definition, right ?

OpenStudy (anonymous):

And the way Lars solves it looks nothing like the way th e video does

ganeshie8 (ganeshie8):

I see..

OpenStudy (anonymous):

Use the limit definition to find the slope of the tangent line to the graph of f at the given point. f(x) = x + 1 ; (3, 2)

OpenStudy (anonymous):

it should show square root of x+1

OpenStudy (anonymous):

:( somebody please help...

ganeshie8 (ganeshie8):

are you using the limit definition I gave u earlier ?

ganeshie8 (ganeshie8):

this one : \[\large f'(x) = \lim \limits_{\Delta x \to 0}~ \dfrac{f(x+\Delta x) - f(x)}{\Delta x}\]

ganeshie8 (ganeshie8):

??

OpenStudy (anonymous):

sorry my laptop died, had to plug in. it looks exactly like that except the F'(x) it's just f(x) and then the rest is the same

ganeshie8 (ganeshie8):

okay good, f'(x) is just a notation used for derivative/slope - forget it for now. lets continue :)

OpenStudy (anonymous):

I end up the last few steps as